undeadknight26 (undeadknight26):

3. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. F(x) = 2x + 1200

undeadknight26 (undeadknight26):

2x + 3y = 1200 i changed it to. F(x) = 2x + 1200

undeadknight26 (undeadknight26):

@amistre64 @HaleyT @agent0smith @charlotte123 @dpasingh @helppls123

OpenStudy (charlotte123):

so the problem is 2x + 3y = 1200 and u wanted to change it to function notation right? ^_^

undeadknight26 (undeadknight26):

yea :D

OpenStudy (charlotte123):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. This is the problem? :)

undeadknight26 (undeadknight26):

yea...

OpenStudy (charlotte123):

this might help: ^_^ http://openstudy.com/study#/updates/52642f68e4b040e33d2c1d6e

undeadknight26 (undeadknight26):

so y = 2?

undeadknight26 (undeadknight26):

@charlotte123 please come back...

OpenStudy (charlotte123):

I'm sorry, someone linked me >.> y = 400, x = 600, and the slope is -2/3 in 2x + 3y = 1200 ^_^

undeadknight26 (undeadknight26):

what was the function notation though?